# Shuffle-R1: Efficient RL framework for Multimodal Large Language Models via Data-centric Dynamic Shuffle

### v4

@@ 4.1 Experimental Setup
We first perform evaluation on in-domain test set of Geometry3K and MMK12. Further, as RL is famous for its strong generalizability, we evaluate our model’s performance on the following representative visual reasoning benchmarks: MathVerse (Zhang et al., 2024b), MathVision (Wang et al., 2024), WeMath (Qiao et al., 2024), MathVista (Lu et al., 2023), HallusionBench (Guan et al., 2024) and ChartQA (Masry et al., 2022). These benchmarks span across math reasoning, visual perception, and chart understanding. We use MathRuler to evaluate questions with free-form ground truths and Gemini-2.0-Flash-001 (Deepmind, 2025) to evaluate questions with multi-choice ground truths. More information in Appendix E.
@@
We use EasyR1 (Yaowei et al., 2025) as our training codebase. We employ Qwen2.5-VL-3B/7B-Instruct (Bai et al., 2025) as base model to verify our method’s generalizability on model scales. Parameters of vision encoder are kept frozen. We set update batch size to 128 and rollout batch size ($G$) to 512. Rollout temperature is set to 1.0 and learning rate is set to 1e61e-6. All experiments are conducted on $8\times$ H800-80G GPUs. For each query, we generate $2N=16$ rollouts. For PTS, we construct $N=8$ pairs and select the top $M=4$ pairs, corresponding to a retention ratio of $\alpha=0.5$, striking a balance between training cost and exploration space. For ABS, we set $T=256$ pairs (512 query-response trajectories) for each sub-sampling batch. We construct shuffled batch by performing $S=8$ rounds of shuffle. Decoding temperature for evaluation is set to 0.5, and we report average pass@1 accuracy of 8 tests to reduce randomness.
@@ 4.2 Main Results
We compare our method with GRPO (Shao et al., 2024) and DAPO (Yu et al., 2025). On Geometry3K (Tab. 2), our 3B model reaches 47.88% accuracy, outperforming GRPO by 5.2% and DAPO by 2.7%.; the 7B model achieves 55.89%, with gains of 3.3% and 1.4%, respectively. On out-of-domain math reasoning tasks, our method further improves average accuracy by 1.96% (3B) and 1.5% (7B) over GRPO, also surpassing DAPO. Consistent gains are observed on HallusionBench and ChartQA, where data distribution diverges more from training. Similar trends hold on the K12 experiments (Tab. 2), with improvements in both in-domain and out-of-domain settings, underscoring the robustness of our framework across tasks and distributions. Compared with the latest GSPO (Zheng et al., 2025), which replaces token-level importance sampling with sequence-level importance sampling, our method also demonstrates superior performance on both in-domain and out-of-domain tasks. All the performance results above indicate our framework’s generalizability on different data distributions and model scales, highlighting the effectiveness of dynamic RL training paradigm. More results in Appendix C.
@@ 4.4 Extension Experimental Results
Model MathVerse MathVision MathVista WeMath HallBench ChartQA Avg.
@@
Model RefCOCO(testA) RefCOCO(testB) RefCOCO+(testA) RefCOCO+(testB) RefCOCOg(test)
@@
Referring Expression Comprehension
@@
Language-only Reasoning
While our initial analysis focused on the RL training dynamics of MLLMs, we conduct an early extension experiment to validate the feasibility of Shuffle-R1 on LLMs. We train Qwen2.5-Math-1.5B/7B-Base (Yang et al., 2024) on Open-S1 dataset (Dang and Ngo, 2025) for 150 steps and evaluate model performance on Math12K (Hendrycks et al., 2021), AIME24, MATH500 (Lightman et al., 2023), GSM8K (Cobbe et al., 2021), GPQA-Diamond (Rein et al., 2023) and OlympiadBench (He et al., 2024). As shown in Tab. 8, our framework delivers significant improvements on all the benchmarks compared to GRPO, demonstrating its potential effectiveness on text-only LLMs.
@@ Acknowledgment
This work was supported by the National Natural Science Foundation of China (NSFC) under Grants No. 62225603 and 623B2038.
@@ Appendix D Experiment Settings
All experiments are conducted on $8\times$ NVIDIA H800-80G-SXM GPUs.
@@ Appendix E Dataset Collection
To fully evaluate model’s reasoning ability, apart from in-domain test set from training dataset, we select several representative benchmarks to examine model’s performance on math reasoning, visual perception, and chart understanding tasks. For math reasoning task, we choose MathVerse (Zhang et al., 2024b), MathVision (Wang et al., 2024), WeMath (Qiao et al., 2024) and MathVista (Lu et al., 2023). We choose HallusionBench (Guan et al., 2024) and ChartQA (Masry et al., 2022) for visual perception and chart understanding tasks respectively. We believe that a good RL algorithm can not only improve in-domain performance, but also have the potential to promote robust generalization to out-of-domain tasks. Examples of evaluation data shown in Fig. 11.

### v5

@@ 4.1 Experimental Setup
We first perform evaluation on in-domain test set of Geometry3K and MMK12. Further, as RL is famous for its strong generalizability, we evaluate our model’s performance on the following representative visual reasoning benchmarks: MathVerse (Zhang et al., 2024b), MathVision (Wang et al., 2024), WeMath (Qiao et al., 2025), MathVista (Lu et al., 2023), HallusionBench (Guan et al., 2024) and ChartQA (Masry et al., 2022). These benchmarks span across math reasoning, visual perception, and chart understanding. We use MathRuler to evaluate questions with free-form ground truths and Gemini-2.0-Flash-001 (Deepmind, 2025) to evaluate questions with multi-choice ground truths. More information in Appendix E.
@@
We use EasyR1 (Yaowei et al., 2025) as our training codebase. We employ Qwen2.5-VL-3B/7B-Instruct (Bai et al., 2025) as base model to verify our method’s generalizability on model scales. Parameters of vision encoder are kept frozen. We set update batch size to 128 and rollout batch size ($G$) to 512. Rollout temperature is set to 1.0 and learning rate is set to 1e61e-6. All experiments are conducted on $8\times$ 80G GPUs. For each query, we generate $2N=16$ rollouts. For PTS, we construct $N=8$ pairs and select the top $M=4$ pairs, corresponding to a retention ratio of $\alpha=0.5$, striking a balance between training cost and exploration space. For ABS, we set $T=256$ pairs (512 query-response trajectories) for each sub-sampling batch. We construct shuffled batch by performing $S=8$ rounds of shuffle. Decoding temperature for evaluation is set to 0.5, and we report average pass@1 accuracy of 8 tests to reduce randomness.
@@ 4.2 Main Results
We compare our method with GRPO (Shao et al., 2024) and DAPO (Yu et al., 2025). On Geometry3K (Tab. 2), our 3B model reaches 47.88% accuracy, outperforming GRPO by 5.2% and DAPO by 2.7%; the 7B model achieves 55.89%, with gains of 3.3% and 1.4%, respectively. On out-of-domain math reasoning tasks, our method further improves average accuracy by 1.96% (3B) and 1.5% (7B) over GRPO, also surpassing DAPO. Consistent gains are observed on HallusionBench and ChartQA, where data distribution diverges more from training. Similar trends hold on the K12 experiments (Tab. 2), with improvements in both in-domain and out-of-domain settings, underscoring the robustness of our framework across tasks and distributions. Compared with the latest GSPO (Zheng et al., 2025), which replaces token-level importance sampling with sequence-level importance sampling, our method also demonstrates superior performance on both in-domain and out-of-domain tasks. All the performance results above indicate our framework’s generalizability on different data distributions and model scales, highlighting the effectiveness of dynamic RL training paradigm. More results in Appendix C.
@@ 4.4 Extension Experimental Results
Setting MathVerse MathVision MathVista WeMath HallBench ChartQA Avg.
@@
Setting RefCOCO(testA) RefCOCO(testB) RefCOCO+(testA) RefCOCO+(testB) RefCOCOg(test)
@@
Referring Expression Comprehension.
@@
Language-only Reasoning.
While our initial analysis focused on the RL training dynamics of MLLMs, we conduct an early extension experiment to validate the feasibility of Shuffle-R1 on LLMs. We train Qwen2.5-Math-1.5B/7B-Base (Yang et al., 2024) on Open-S1 dataset (Dang and Ngo, 2025) for 150 steps and evaluate model performance on Math12K (Hendrycks et al., 2021), AIME24, MATH500 (Lightman et al., 2023), GSM8K (Cobbe et al., 2021), GPQA-Diamond (Rein et al., 2024) and OlympiadBench (He et al., 2024). As shown in Tab. 8, our framework delivers significant improvements on all the benchmarks compared to GRPO, demonstrating its potential effectiveness on text-only LLMs.
@@ Acknowledgment
This work was supported by the National Natural Science Foundation of China (NSFC) under Grants No. 62225603, 62576147 and 623B2038.
@@ Appendix D Experiment Settings
All experiments are conducted on $8\times$ 80G GPUs.
@@ Appendix E Dataset Collection
To fully evaluate model’s reasoning ability, apart from in-domain test set from training dataset, we select several representative benchmarks to examine model’s performance on math reasoning, visual perception, and chart understanding tasks. For math reasoning task, we choose MathVerse (Zhang et al., 2024b), MathVision (Wang et al., 2024), WeMath (Qiao et al., 2025) and MathVista (Lu et al., 2023). We choose HallusionBench (Guan et al., 2024) and ChartQA (Masry et al., 2022) for visual perception and chart understanding tasks respectively. We believe that a good RL algorithm can not only improve in-domain performance, but also have the potential to promote robust generalization to out-of-domain tasks. Examples of evaluation data shown in Fig. 11.